# SoK: Consensus for Fair Message Ordering

### v1

@@ III-A2 FIFO via Timestamping
Based on the existing proposals, there are two ways to timestamp a message. The first way is to make Byzantine agreements using the median value of the timestamps that the nodes locally receive of a message [29, 30, 31, 32, 34, 35, 38]. The median value is used because it eliminates the effect of the malicious timestamp proposals from the Byzantine nodes. The median timestamp is always within the range of time the message was received by the honest nodes.
Given a set of messages $S$, let $T^{S}_{i}$ denote the list of timestamps at which every message of $S$ is locally observed by node $i$. FIFO ordering via median timestamping is defined as the following ordering rule:
@@
Table I provides an example scenario of FIFO via median timestamping. Based on the decided timestamp, message 1 is ordered before message 2. However, a problem with using timestamps from a single node (the median) is that the node can manipulate the relative ordering of transactions. If node 2 is malicious, it can change the order of the messages simply by sharing the flipped local timestamps for receiving the two messages. Although it is believed that such manipulation is difficult in practice for an adversary [29, 30] as it requires an adversary to submit a timestamp that is exactly the median, the possibility of such manipulation cannot be fully eliminated.
TABLE I: An example of taking median timestamp to order messages
Message Node 1 Node 2 Node 3 Decided timestamp
@@
When involving some median timestamp $t_{m}$, it always comes from one single node (or two nodes) in the network. Whenever $t_{m}$ is used to compare with timestamps from any other nodes in the network, there must be an assumption that these two local clocks are synchronized, otherwise the comparison does not provide any conclusions. Since all existing median timestamping proposals involve cross-node timestamp comparisons, they all require synchronized clocks. This does not mean median timestamping methods are infeasible, as BFT clock synchronization protocol exists and it can be applied to ensure synchronized clocks [43].
Byzantine agreement of the median value given a set of values has been extensively studied [44], and it is also used in this line of research to find out the best median value among the honest nodes [38]. However, we argue that the problem here is somewhat different from finding a median value. Instead of finding the median value among honest proposals, the aim of this problem is to find the minimum honest value, as it is the closest timestamp to when the message is actually sent. This brings up a research gap that has not yet been studied, that is, how accurate the timestamps derived from these protocols are from the actual time when the message is sent by the sender, and how to minimize the deviation between the actual and derived timestamps.
@@
In this category of approach, both proposals [36, 37] assume a message is broadcast from one node to all other nodes. In the case that the network is not synchronized and the network delay fluctuates, nodes can not reach an agreement because the resulting timestamping is not within the acceptable range of their expectation, therefore liveness is lost. Given such strict conditions, the scalability and resilience of this approach is yet to be verified.
III-B Random Ordering Rule
Another idea of fair ordering is to randomly order messages [45, 46, 47, 48, 49, 50, 51].
Given a set of messages $S$, denote $R$ as the randomness agreed by all nodes. Random ordering is defined as:
@@ III-B Random Ordering Rule
The core of random ordering is to obtain a randomness that all nodes agree. The existing proposals [45, 46, 47, 48, 49, 50, 51] use information from the previous blocks as the public randomness. To convert randomness into a message selection criteria, a hash function is used to combine the randomness and the messages to output a pseudo-random value for each message at each message selection round, then the messages with the lowest pseudo-random values are selected.
@@
Early proposals for random ordering [45, 46, 47, 48] are non-deterministic because these proposals do not require a unified set of messages $S$ as the input of the ordering rules. A proposer derives a randomly ordered message set using its own message pool, and the peers verify the proposal by inferring the probability of the proposer being honest given their own message pools. For example, if too many messages in the proposal are not known by a peer, this peer probabilistically rejects the proposal. The order derived from these proposals, even though agreed by all nodes, may still contain maliciously constructed messages from the proposer to fit in a designated order position. Recent proposals propose to inject randomness by randomly permute messages on a set of messages the nodes agreed to commit [49, 50, 51], providing a deterministic way of enforcing randomness.
III-C Blind Ordering Restriction
@@
Another line of research is blind ordering [52, 53, 45, 46, 54, 55, 56, 57, 37, 58, 59, 60, 61, 62, 63, 64, 65, 66, 62, 63, 64]. These proposals use threshold cryptography [67, 68] to encrypt the messages before they are committed so that nodes are unable to read the content of the messages when ordering. It requires message senders to perform encryption that allows only a threshold of the nodes (e.g. three out of four) to collectively reveal the messages after they are committed. Except for threshold cryptography, computational time lock puzzles can also be used to keep messages encrypted before committing [69]. However, the time lock puzzles have the intrinsic disadvantage of having unstable puzzle-solving time [70], meaning that messages can be revealed earlier than expected. Therefore, encrypting with time lock puzzles is less secure.
@@
Strictly speaking, this approach on its own is not an ordering rule as it still allows nodes to arbitrarily order messages, therefore it is titled as a restriction other than a rule. It can be orthogonally applied along with random ordering [45, 46, 47, 48, 50, 51] or FIFO ordering [37, 38, 33].
@@
IV Fair Consensus Factory: Adding Fairness into Consensus Protocols
To the fair ordering rules in a distributed ledger, the nodes need to reach an agreement on the order results, namely, to embed fair ordering rules into the consensus protocols. This section presents the fair consensus factory, a framework to add order fairness to any byzantine fault-tolerant committee-based consensus protocols, which are commonly used in distributed ledgers.
IV-A Selecting Fair Ordering Rules
@@
Before bringing a consensus into the factory, a decision to make is to select which ordering rules to apply. Figure 4 shows the fair ordering options a consensus protocol has. A consensus can either use random ordering, FIFO via ranked voting, or FIFO via timestamping, with an optional add-on of requiring the nodes to apply the ordering rule without knowing their content. Details of these options are covered in the previous section.
@@
Figure 4: Fair consensus factory menu: select ordering rules
IV-B Process of Adding Fairness
The idea of the process is to, first, synchronize the local states of the nodes including the local message pools and the required metadata. Subsequently, all nodes can execute the ordering rule to derive the ordered message set of this round that they should commit. This is not the only possible workflow for fair ordering consensus. There are also proposals [55, 10, 71, 72] to use TEE message pools to enforce transaction selection rules, so that only the leader of that consensus round needs to execute the rule and others can trust the result is correct. However, this creates centralization risks to trust the hardware manufacturer [73], which is not desired in the decentralized setting. Another potential way that only requires one node to execute the ordering rule is to create zero-knowledge proof for the peers to verify whether the computation is correctly executed [1]. However, this remains a future research direction as it creates extra heavy computation for the leader and, to the best of our knowledge, there are not yet any evaluations showing that this approach can reduce latency. With leader performance bottleneck as a known problem for leader-based consensus [74], this step of running additional computations is expected to contribute significantly to higher latency.
Therefore, the fair consensus factory sticks to the approach of all-to-all synchronizing, and all nodes execute rules locally. It takes three steps to modify a normal consensus protocol to achieve so:
Uninstall the Propose Stage
Without an enforced ordering rule, normal consensus protocols allow nodes to freely propose an ordered message set at a propose stage. This stage is not needed in a fair ordering consensus because the message set to commit should be derived deterministically from the ordering rules. Therefore, the first step of the process is to remove the propose stage.
Install the Synchronization Stage
Normal consensus can be executed with asynchronized message pools for each node. However, if message pools are not synchronized in fair ordering consensus different local views of the message pool greatly increase the opportunity of getting unfair results (as shown in almost all existing proposals) and may be prone to attacks [27]. Therefore, a synchronization stage is required for nodes to update each other their local message pools before ordering. If the consensus is to be processed into a FIFO ordering consensus, this synchronization stage also involves exchanging the local metadata (detailed in the previous section) required to execute the ordering rule.
Note that in FIFO ordering, nodes are expected to not only synchronize in this synchronization stage of a consensus process, but it is also required that they constantly exchange local messages at the network layer before starting a consensus process. This is to populate the metadata (e.g. timestamps of locally receiving the messages for FIFO via timestamping rules) to be used in the consensus process.
Install the Rule Execution Stage
The purpose of the above state synchronization stage is to allow all nodes to execute the fair ordering rule with the same input message set and metadata. With the synchronized states, they can locally execute the ordering rule and output the same ordered message set, which is the state they need to reach an agreement upon and commit to in this consensus round. Unlike the synchronization stage which requires all-to-all communication to exchange local states, this stage requires only local computations for each node. After this stage, normal consensus stages can be applied to reach an agreement, commonly referred to as the pre-commit and commit phase.
IV-C Different Processing Approaches for Leader-based and Leaderless Consensus
Leader-based [29, 59, 1] (to name a few, majority of consensus used in distributed ledgers are leader-based) and leaderless [75, 11, 76, 77, 78, 20] are two categories of committee-based consensus protocol. This categorization makes a difference in fairness consensus factory because it requires different approaches to processing them.

### v3

@@ 3.1.2 FIFO via Timestamping
Based on the existing proposals, there are two ways to timestamp a message. The first way is to make Byzantine agreements using the median value of the timestamps that the nodes locally receive for a message [27, 28, 29, 30, 32, 33, 36]. The median value is used because it eliminates the effect of the malicious timestamp proposals from the Byzantine nodes. The median timestamp is always within the range of time the message was received by the honest nodes.
Given a set of messages $S$, let $T^{S}_{i}$ denote the map of timestamps at which every message of $S$ is locally observed by node $i$. FIFO ordering via median timestamping is defined as the following ordering rule:
@@
Using the median timestamp $t_{m}$ implies that it comes from one single node (or two nodes) in the network. Table I provides an example scenario of FIFO via median timestamping. Based on the decided timestamp, message 1 is ordered before message 2. However, a problem with using timestamps from a single node (the median) is that the node can manipulate the relative ordering of messages. If node 2 is malicious, it can change the order of the messages simply by claiming message 1 is received at 3:03, and message 2 is received at 3:02. Although it is believed that such manipulation is difficult in practice for an adversary [27, 28] as it requires an adversary to submit a timestamp that is exactly the median, the possibility of such manipulation cannot be fully eliminated.
TABLE I: An Example of Taking the Median Timestamp to Order Messages
Message Node 1 Node 2 Node 3 Output
@@
Existing definitions [28, 27, 1] attempting to formalize the fairness guarantee provided by the median timestamping approach do not precisely cover the above scenario. Kelkar et al. [1] conclude that, in this approach, if all honest nodes receive a message $m_{1}$ before any honest node receives another message $m_{2}$, then all honest nodes output $m_{1}$ before $m_{2}$. However, this is not the case if a malicious node swaps the median timestamps of the two transactions. Therefore, we provide a definition that captures the fairness guarantee provided by this approach:
Definition 2 ($f$-Robust Median Order)
Let $n$ be the number of nodes and $f$ the maximum number of Byzantine nodes. Let $\mathcal{T}(a)$ and $\mathcal{T}(b)$ be the multisets of timestamps for transactions $a$ and $b$, respectively, with $n$ entries each. Let $\mathrm{ts}(a)$ denote the median of $\mathcal{T}(a)$.
We say that the relative order $\mathrm{ts}(a)<\mathrm{ts}(b)$ is $f$-robust if:
$T_{(n/2+f+1)}(a)<T_{(n/2-f)}(b)$
where $T_{(i)}(a)$ denotes the $i$-th smallest timestamp in $\mathcal{T}(a)$, and likewise for $b$.
This definition accounts for an adversary controlling up to $f$ timestamps per transaction that can reverse the order by manipulating the median, and only guarantees fair ordering between messages that are not vulnerable to such an attack, given that they share fewer overlapping timestamps.
Moreover, whenever $t_{m}$ is used to compare with timestamps from any other nodes in the network, there must be an assumption that these two local clocks are synchronized, otherwise the comparison does not provide any conclusions. Since all existing median timestamping proposals involve cross-node timestamp comparisons, they all require synchronized clocks. This does not mean median timestamping methods are infeasible, as BFT clock synchronization protocol exists and it can be applied to ensure synchronized clocks [41, 42].
Byzantine agreement of the median value given a set of values has been extensively studied [43], and it is also used in this line of research to find out the best median value among the honest nodes [36]. However, we argue that the problem here is somewhat different from finding a median value. Instead, the aim of this problem is to find the minimum honest value, as it is the closest timestamp to when the message is actually sent. This brings up a research gap that has not yet been studied, that is, how accurate the timestamps derived from these protocols are from the actual time when the message is sent by the sender, and how to minimize the deviation between the actual and derived timestamps.
@@
In this category of approach, both proposals [34, 35] assume a message is broadcast from one node to all other nodes. In the case of unstable network delay and the network is not synchronized, nodes can not reach an agreement because the resulting timestamping is not within the acceptable range of their expectation, therefore liveness is lost. Given such strict conditions, the scalability and resilience of this approach are yet to be verified.
3.1.3 Comparison of FIFO Protocols
A comparison of FIFO protocols can be found in Themis [1]. We provide a more comprehensive version of such a comparison in Table II. The table only includes proposals from which clear, fair ordering algorithms with fairness guarantees are proposed.
TABLE II: Comparison of FIFO Ordering Consensus Protocols Across Design, Assumptions, and Performance Dimensions
Design Security Assumptions Performance
Proposal Leader-Based/Leaderless Ordering Rule Ordering Guarantee Byzantine Fault Tolerance Network Model Sync. Clocks Comm. Complexity Ordering Complexity
Aequitas [19] Leaderless Voting $\gamma$-batch order fairness $n\geq\frac{2f(\gamma+1)}{2\gamma-1}$ (weak liveness)a Asynchronous No $O(n^{3})$ $O(m^{2}\cdot n)$
Quick fairness [21] Leaderless Voting $\gamma$-batch order fairness 0b Asynchronous No $O(n^{2})$ $O(m^{2}\cdot n)$
Themis [1] Leader-based Voting $\gamma$-batch order fairness $n\geq\frac{2f(\gamma+1)}{2\gamma-1}$ Partially synchronous No $O(n^{2})$ $O(m^{2}\cdot n)$
Pompē [27] Leader-based Timestamping $f$-Robust Median Order $n\geq 3f+1$ Palatially synchronous Yes $O(n^{2})$ $O(n\log n)$
Wendy [28] Leader-basedc Timestamping $f$-Robust Median Order $n\geq 3f+1$ Asynchronousd Yes $O(n^{2})$ $O(n\log n)$
Hashgraph [30] Leaderless Timestamping $f$-Robust Median Order $n\geq 3f+1$ Asynchronous Yes $O(n)$ $O(n\log n)$
  • Note: $n$: the number of nodes in the network; $f$: the number of Byzantine faults; $m$: the number of messages in a round of consensus.
  • a
    In the case of a long Condorcet cycle, Aequitas does not provide a liveness guarantee even when all nodes are honest.
  • b
    The proposal loses liveness in the presence of any Byzantine nodes.
  • c
    Wendy is only a pre-protocol for a consensus protocol, but this pre-protocol itself requires a leader.
  • d
    Also depends on the specific consensus protocol Wendy integrates with.
Proposals based on FIFO via voting incur higher computational complexity in their ordering algorithms. This is primarily due to the need to construct a dependency graph of messages, based on the local orderings reported by all participating nodes. In contrast, timestamping-based approaches tend to have lower ordering complexity, as their primary computational bottleneck lies in sorting timestamps.
It is also important to note that voting-based proposals generally tolerate fewer Byzantine faults than their timestamping counterparts, even under the most relaxed fairness parameter (i.e., when $\gamma=1$). Both Aequitas [19] and Themis [1] can tolerate up to one-fourth of the nodes being Byzantine. This limitation arises from the fairness condition, which requires that at least $\gamma n$ nodes observe one message before another for the ordering to be considered fair. Since up to $f$ of these nodes may behave maliciously and lie about their observations, the protocol must ensure that $\gamma n-f>\frac{n}{2}$ to maintain correctness, thus limiting the fault tolerance.
3.2 Random Ordering Rule
3.2.1 Methods for Random Ordering
Another approach of fair ordering is to randomly order messages [44, 45, 46, 47, 48, 49, 50].
Given a set of messages $S$, denote $R$ as the randomness agreed upon by all nodes. Random ordering is defined as:
@@ 3.2.1 Methods for Random Ordering
The core of random ordering is to obtain a randomness that all nodes agree on. The existing proposals [44, 45, 46, 47, 48, 49, 50] use information from the previous blocks as the public randomness. To convert randomness into a message selection criterion, a hash function is used to combine the randomness and the messages to output a pseudo-random value for each message at each message selection round. Then the messages with the lowest pseudo-random values are selected.
@@
Early proposals for random ordering [44, 45, 46, 47] are non-deterministic because these proposals do not require a unified set of messages $S$ as the input of the ordering rules. A proposer derives a randomly ordered message set using its own message pool, and the peers verify the proposal by inferring the probability of the proposer being honest given their own message pools. For example, if too many messages in the proposal are not known by a peer, this peer probabilistically rejects the proposal. The order derived from these proposals, even though agreed by all nodes, may still contain maliciously constructed messages with specific hashes to fit in a designated order position. With this approach, no deterministic fair ordering guarantee can be formulated based on this approach.
To prevent so, recent work proposes to randomly permute a set of committed messages for execution [48, 49, 50], providing a deterministic way of enforcing randomness. We provide a formal definition of the ordering guarantee produced by such a permutation approach.
Definition 3 (Permuted Ordering)
Let $M=\{m_{1},m_{2},\dots,m_{n}\}$ be a set of $n$ committed messages, and let $\sigma:M\to\{1,\dots,n\}$ be a random bijection drawn uniformly at random from the set of all permutations over $M$. The protocol enforces a permuted ordering if for all distinct messages $a,b\in M$, the following holds:
$\Pr[\sigma(a)<\sigma(b)]=\Pr[\sigma(b)<\sigma(a)]=\frac{1}{2}.$
3.2.2 Comparison of Random Ordering Protocols
TABLE III: Comparison of Random Ordering Consensus Protocols Across Design, Assumptions, and Performance Dimensions
Design Security Assumptions Performance
Proposal Leader-Based/Leaderless Ordering Rule Ordering Guarantee Byzantine Fault Tolerance Network Model Sync. Clocks Comm. Complexity Ordering Complexity
Helix [44] Leader-based Random Selection Probabilistic random selection $n\geq 3f+1$ Synchronous No $O(n^{2})$ $O(m\log m)$
$\Pi^{3}$ [48]a Leader-based Permutation Permuted Ordering N/A N/A No N/A $O(m)$
MEVade [50]a Leader-based Permutation Permuted Ordering N/A N/A No N/A $O(m)$
BlindPerm [49] Leader-based Permutation Permuted Ordering $n\geq 3f+1$ Partially synchronous No $O(n^{2})$ $O(m)$
  • Note: $n$: the number of nodes in the network; $f$: the number of Byzantine faults; $m$: the number of messages in a round of consensus.
  • a
    $\Pi^{3}$ and MEVade are modules that can be adapted to different leader-based protocols, not an actual consensus protocol.
Table III lists four random ordering proposals. The permutation-based approach offers the lowest ordering complexity among fair ordering protocols. It avoids the overhead of constructing dependency graphs or sorting timestamps, making it computationally lightweight. Additionally, it does not rely on synchronized clocks. Prior work [48, 50] has also demonstrated that permutation schemes can be easily integrated with a variety of consensus protocols, highlighting their modularity and practicality. These properties make the permutation approach a lightweight solution against MEV.
3.3 Blind Ordering Restriction
@@
Another line of research is blind ordering [51, 52, 44, 45, 53, 54, 55, 56, 35, 57, 58, 59, 60, 61, 62, 63, 64, 65, 61, 62, 63]. These proposals use threshold cryptography [66, 67] to encrypt the messages before they are committed so that nodes are unable to read the content of the messages when ordering. It requires message senders to perform encryption that allows only a threshold of the nodes (e.g. three out of four) to collectively reveal the messages after they are committed. Except for threshold cryptography, computational time lock puzzles can also be used to keep messages encrypted before committing [68]. However, the time lock puzzles have the intrinsic disadvantage of having unstable puzzle-solving time [69], meaning that messages can be revealed before ordering. Therefore, encrypting with time lock puzzles is less secure.
@@
Strictly speaking, this approach on its own is not an ordering rule as it still allows nodes to arbitrarily order messages, therefore, it is considered rather a restriction than a rule. It can be orthogonally applied along with random ordering [44, 45, 46, 47, 49, 50] or FIFO ordering [35, 36, 31].
@@
4 Fair Consensus Factory: Building Efficient Fair Ordering Consensus Protocols
Inspired by the existing proposals that decouple fair ordering algorithms with consensus protocols [28, 22, 48, 50], the design of fair ordering consensus protocols can be improved by treating fair ordering as a modular refinement to existing consensus systems. We introduce the Fair Consensus Factory, a framework that systematically transforms Byzantine fault-tolerant consensus protocols to enforce fair message ordering. This section formalizes this idea as a protocol transformation, outlines the design space of ordering rules, and presents a transformation-based method for adding order fairness into both leader-based and leaderless consensus protocols.
4.1 Adding Order Fairness to Consensus Protocol
Let $\mathcal{P}$ denote the class of Byzantine fault-tolerant consensus protocols that satisfy standard safety and liveness properties, but do not satisfy any order fairness property in Definition 1, 2, 3. Each protocol $P\in\mathcal{P}$ assumes a distributed set of nodes, a message dissemination layer, and a decision procedure that determines a sequence of committed transactions.
Given a deterministic fair ordering rule $r$, the Fair Consensus Factory transforms $P$ into $P_{\text{fair}}$, such that $P_{\text{fair}}$ is a protocol in which message ordering adheres to a deterministic, fair ordering function $r$ shared by all honest nodes.
This transformation is realized through two transformation rules applied to $P$:
Transformation Rule 1: State Synchronization
In standard consensus protocols, nodes may maintain inconsistent views of pending messages due to asynchronous delivery, selective dissemination, or adversarial interference. To enforce fairness, we introduce a synchronization step that ensures all honest nodes agree on the input set and associated metadata prior to ordering. Specifically, honest nodes must establish a common view of the candidate message set $S$ and its metadata $D$ through explicit communication or implicit dissemination during consensus rounds. This shared state is a precondition for fair ordering, and prevents adversaries from biasing the outcome by withholding or selectively propagating messages.
Transformation Rule 2: Local Rule Execution
Once a synchronized view of $S$ and $D$ is established, each honest node independently applies the fair ordering function $r$ to produce an ordered output sequence. Because $r$ is deterministic and operates on agreed-upon inputs, all honest nodes produce the same result without relying on a centralized proposal or aggregation phase. This eliminates opportunities for order manipulation and aligns the protocol output with a predefined fairness criterion.
The protocol no longer relies on arbitrary leader proposals or aggregation heuristics to determine message order. Instead, fairness is achieved by synchronizing input views and applying a locally computable, globally consistent rule.
The transformation only changes how $P$ derives a proposal to agree on, but it does not change how $P$ reaches agreement on the proposal.
4.2 Choosing Ordering Rules
@@
The factory supports a variety of fair ordering rules $r$, selected depending on desired fairness criteria and system assumptions. Figure 4 shows options for ordering rules. A consensus can either use random ordering, FIFO via ranked voting, or FIFO via timestamping, with an optional add-on of requiring the nodes to apply the ordering rule without knowing their content. Details of these options are covered in the previous section.
@@
Figure 4: Fair Consensus Factory menu: selection of ordering rules
4.3 Application to Leader-Based and Leaderless Protocols
The Fair Consensus Factory applies to both leader-based and leaderless protocols by specializing in how transformation rules are embedded into the protocol logic. Figure 5 compares the transformation applied to each protocol family.
Leader-Based Consensus
A leader-based consensus protocol (e.g., HotStuff [70], Alea [58]) proceeds through the following phases per round:
  1. 1.
    Proposal: A designated leader collects a set of pending transactions $S$ from its local pool and produces an ordered sequence $O$ from an arbitrary ordering policy.
  2. 2.
    Agreement: The leader broadcasts the proposal $O$ to other nodes. If they validate it, they vote to accept it, and the protocol reaches consensus on $O$.
The Fair Consensus Factory transforms this process as follows:
  • At the beginning of the consensus round (e.g., during leader election) or at the last phase of the last consensus round, all nodes exchange $S_{i}$ and $D_{i}$ to synchronize the input set $S$ and associated metadata $D$.
  • Once $S$ and $D$ are agreed upon, each honest node independently computes $O=r(S,D)$ using the fair, deterministic ordering rule $r$.
  • The result $O$ is passed into the standard voting and agreement phases of the protocol, unmodified.
For leader-based consensus, state synchronization can often be amortized or embedded in leader election or finalization messages of the previous consensus round, minimizing latency overhead.
Leaderless Consensus
Leaderless consensus protocols (e.g., Narwhal/Bullshark [71], Avalanche [10], Mysticeti [72]) typically follow a different structure:
  1. 1.
    Exchange proposals: Each node $i$ proposes a local input set $S_{i}$ of messages.
  2. 2.
    Aggregation: The protocol aggregates proposals $S_{1},...,S_{n}$ into a combined set $S$, and applies an predefined ordering policy to derive a sequence $O$.
  3. 3.
    Agreement: Nodes reach consensus on $O$ using the standard agreement procedure of the protocol.
This process already satisfies the communication requirements of the transformation rules, therefore, the Fair Consensus Factory only modifies message exchange and the ordering rule as follows:
  • The exchange proposal phase is also used to exchange metadata $D_{i}$. As a result, honest nodes synchronize on a common input set $S$ and metadata $D$.
  • The originally defined aggregation rule is replaced by the fair ordering rule $r$. Each node locally computes $O=r(S,D)$.
  • The agreed-upon output $O$ is then passed into the standard agreement procedure of the protocol.